OpenStudy (anonymous):

Solve exactly pi^(-2x+1)=e^x

OpenStudy (jdoe0001):

\(\bf \pi^{-2x+1}=e^x\qquad \textit{taking "ln" to both sides}\\ \quad \\ log_e(\pi^{-2x+1})=log_e(e^x)\\ \quad \\ \textit{log cancellation rule of }\quad log_aa^x=x\qquad thus\\ \quad \\ log_e(\pi^{-2x+1})=log_e(e^x)\implies log_e(\pi^{-2x+1})=x\) ... so... what do you think about the left side? can we ... do something with the exponent? maybe a log exponent rule?

OpenStudy (jdoe0001):

\(\Large \bf \textit{recall that}\quad log_a(x^y)\implies ylog_ax\)

OpenStudy (jdoe0001):

... so... what do you think?

OpenStudy (isaiah.feynman):

@jdoe0001 is always there to solve problems I find interesting. :P

OpenStudy (anonymous):

Thank you! So now, can we change it into exponential form, so that, e^x=-2x+1, or does that not help?

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

\(\bf \pi^{-2x+1}=e^x\qquad \textit{taking "ln" to both sides}\\ \quad \\ log_e(\pi^{-2x+1})=log_e(e^x)\implies log_e(\pi^{-2x+1})=x\\ \quad \\ (-2x+1)log_e(\pi)=x\) so if you distribute the left-side,... what would it look like?

OpenStudy (jdoe0001):

\(\bf \pi^{-2x+1}=e^x\qquad \textit{taking "ln" to both sides}\\ \quad \\ log_e(\pi^{-2x+1})=log_e(e^x)\implies log_e(\pi^{-2x+1})=x\\ \quad \\ (-2x+1)log_e(\pi)=x\) so if you distribute the left-side,... what would it look like?

OpenStudy (anonymous):

What is it that we are distributing here?

OpenStudy (jdoe0001):

\(\bf {\Large (-2x+1)log_e(\pi)}=x\)

OpenStudy (anonymous):

So if I plug in ln for loge I would have (-2x+1)lnpi=x ? Which would then be ln(-2x)+ln+lnpi ?

OpenStudy (jdoe0001):

well... you can think of it as say .... \(\bf (-2x+1)log_e(\pi)=x\qquad log_e(\pi)=a\implies (-2x+1)a=x\) if you distribute that... what would it give you?

OpenStudy (anonymous):

Oh okay, -2xa+a=x

OpenStudy (jdoe0001):

yeap.. so \(\bf (-2x+1)log_e(\pi)=x\qquad log_e(\pi)=a\implies (-2x+1)a=x\\ \quad \\ \implies -2xa+a\implies -2xlog_e(\pi)+log_e(\pi)\qquad thus\\ \quad \\ (-2x+1)log_e(\pi)=x\implies (-2x)log_e(\pi)+log_e(\pi)=x\\ \quad \\ (-2x)log_e(\pi)+log_e(\pi)-x=0\implies (-2x)log_e(\pi)-x=-log_e(\pi)\\ \quad \\ \textit{now taking common factor to the left-side}\\ \quad \\ x[]=\implies x=\cfrac{-log_e(\pi)}{-2log_e(\pi)-1}\)

OpenStudy (jdoe0001):

hmm hhehhe.l missing a piece.... there... ok

OpenStudy (jdoe0001):

\(\bf (-2x+1)log_e(\pi)=x\qquad log_e(\pi)=a\implies (-2x+1)a=x\\ \quad \\ \implies -2xa+a\implies -2xlog_e(\pi)+log_e(\pi)\qquad thus\\ \quad \\ (-2x+1)log_e(\pi)=x\implies (-2x)log_e(\pi)+log_e(\pi)=x\\ \quad \\ (-2x)log_e(\pi)+log_e(\pi)-x=0\implies (-2x)log_e(\pi)-x=-log_e(\pi)\\ \quad \\ \textit{now taking common factor to the left-side}\\ \quad \\ x[-2log_e(\pi)-1]=-log_e(\pi)\implies x=\cfrac{-log_e(\pi)}{-2log_e(\pi)-1}\)

OpenStudy (anonymous):

Great! So then, is it acceptable to cancel out the negative signs which would result in ln(pi)/(1+2ln(pi))...?

OpenStudy (jdoe0001):

sure, yes

OpenStudy (anonymous):

\[\ln(\pi)/ (1+2\ln(\pi))\]

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Great, thank you so much!

OpenStudy (jdoe0001):

yw